Summary for owl shift; Nov. 24, 1999

Angela Bellavance (bellavan@rice.edu)
Wed, 24 Nov 1999 07:04:48 -0600 (CST)

Shift Summary for owl shift, 24/Nov/99

On Shift: A. Alavi-Harati, A. Bellavance

Hours of beam delivered at all intensities: 6.5
Hours of beam delivered at nominal intensity: 6.5
Hours of E799 data (special runs not included): 6.3
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: energy conservations and TeV aborts,
starting a run

========================================================================
Runs ended this shift:
Run 15040 - ?M events - BEAM_799 (ongoing)

Tapes brought to FCC this shift: none
(including tapes for online split)
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: Fairly smooth E799 datataking. DAQ complains if the
SEM gets above 1e13. Beam delivery is pretty good but has an occasional
insensity "spike" slightly above 1e13 (can be seen on the spill scope).

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
00:00 Muon run finishing, getting ready to take first
real data since 24hr shutdown.
15040 1 00:15 Start BEAM_799 run.
" 6-15 00:23-34 Energy conservation period.
" 21-27 00:40-51 Energy conserv.
" 31-40 00:55-1:08 Energy conserv.
" 69 01:47 TeV abort for 1 spill.
" 142 03:26 TeV abort for 1 spill.
" 151 03:38 TeV abort for 1 spill.
" 166 03:58 TeV abort for 1 spill.
04:10 Start getting occasional "HCC inefficient"
warnings for CsI index #166.
" 183 04:20 TeV abort, beam gone.
" 187 04:26 Beam returns.
05:31-5 Energy conservation period.
05:55 Ashkan somehow manages to kill the window manager
of the "DISPLAY" terminal. Had to power cycle
that terminal, so online kdisp display is gone
until next run.
06:55 Tape UPK024 mounted on ktev1, tps3d2.
" 301 06:58 Energy conservation period (1 spill).

========================================================================
Follow up Notes on Previous Shift

CsI constants (and Xts) are updated but not yet loaded because we are
waiting for the beam to get crappy / go away for a while to stop and
re-start the run.

========================================================================
Items to Follow up on Next Shift

- keep an eye on CsI index #166 (email already sent to Julie).
- the onsplit monitoring window has disappeared (but the process itself
is still running). Anyone know how to get the window re-started? (The
mount-request window is still there and working.)
- need to start a new gas checklist sheet. (blanks are where?)
- update CsI constants and Xts when there is a break in running.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke